1. A particle is moving in a straight line with initial velocity u and retardation av, where v is the velocity at any time t [a] the particle will cover a total distance u/a [b] the particle will come to rest after 1/a [c] the particle will continue to move for a very long time [d] the velocity of the particle will become u/2 after time 1/a

Grade:11

2 Answers

Sandeep Bhamoo
5 years ago
Here acceleration = dv/dt = av, hence on integrating, Ln(u/v)= at now looking to third option when final velocity is zero Ln(u/0)=infinity =a•t ,means at that moment time would be infinity means the body comes to rest after a long period of time. Please do approve answer if correct
Khimraj
5 years ago
-dv/dt = av
-dv/v = adt
\int_{u}^{v}dv/v= -a\int dt
ln(v/u) = -at
So the particle will continue to move for a very long time.